OpenStudy (anonymous):

Differentiate the function: G(u)= ln sqrt((3u+6)/(3u-6)) The answer is -2/(u^2 -4) but I am not sure how to get to that answer. Steps with explanations would be appreciated!

OpenStudy (anonymous):

\[G(u) =\sqrt{\frac{3u+6}{3u-6}}\]

OpenStudy (anonymous):

you can either use the product rule or the quotient rule to solve this problem...

OpenStudy (anonymous):

\[G(u)=\ln \sqrt{\frac{ 3u+6 }{ 3u-6 }}\] is the equation. I am solving for G'(u)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ok i will move with u step by step

OpenStudy (anonymous):

STEP ONE SIMPLIFY THE LOG

OpenStudy (anonymous):

\[\ln(3u+6)^{1/2}-\ln(3u-6)^{1/2}\] Thats as far as I got

OpenStudy (anonymous):

simplify it even further what happens to the power?

OpenStudy (anonymous):

oh, it goes in front, duh

OpenStudy (anonymous):

I forget what to do from here though. ln is not a natural number, so i cant use chain rule. \[\frac{1}{2}\ln(3+6)-\frac{1}{2}\ln(3-6)\]

OpenStudy (anonymous):

ok what is the diff of 1/2ln(x)?

OpenStudy (anonymous):

well u can use the chaing rule

OpenStudy (anonymous):

d(ln(x)/dx=1/x

OpenStudy (anonymous):

ah, I missed that. I must have not written it in my notes.

OpenStudy (anonymous):

so????

OpenStudy (anonymous):

so i use chain rule for both the 1/2ln(3u+6) and 1/2ln(3u-6) separately? Im stuck trying to figure out how to deal with the ln

OpenStudy (anonymous):

is it just (1/2)(1/3u+6) - (1/2)(1/3u-6)?

OpenStudy (anonymous):

ok i will show u

OpenStudy (anonymous):

can u see the 3 next to the u

OpenStudy (anonymous):

or is it \[[\frac{ 1 }{ 2 }(\frac{ 1 }{ 3u+6 })+\ln(3u+6)] -[\frac{ 1 }{ 2 }(\frac{ 1 }{ 3u-6 })+\ln(3u-6)]\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

look carefully

OpenStudy (anonymous):

i will solve it for u

OpenStudy (anonymous):

you can factor it out, but is it unnecessary?

OpenStudy (anonymous):

the last bit you did was right (1/2)(1/3u+6) - (1/2)(1/3u-6)

OpenStudy (anonymous):

now the right thing is (1/2)(3/3u+6) - (3/2)(1/3u-6)

OpenStudy (anonymous):

can u see the 3s

OpenStudy (anonymous):

yes, where did you get them from? did you just put in (1/3) and (3/1) to keep it balanced?

OpenStudy (anonymous):

from my pocket hehehe, well by using the chain rule

OpenStudy (anonymous):

look let x = 3u+6, dx =3du agree

OpenStudy (anonymous):

ln(u) -diff-> 1/u

OpenStudy (anonymous):

right, i get ln(u) => 1/u, but what about the dx=3du?

OpenStudy (anonymous):

oh, you are just saying that dx where x=3u+6 is equal to 3

OpenStudy (anonymous):

and that is where the 3 comes from because of chain rule!

OpenStudy (anonymous):

the cahin rule\[\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}\]

OpenStudy (anonymous):

have u seen this rule

OpenStudy (anonymous):

not written like that, can you express it in F(x) and G(x), or u and v?

OpenStudy (anonymous):

i have in my notes [f(x)/g(x)]' = (f'(x)g(x)-g'(x)f(x))/g(x)^2

OpenStudy (anonymous):

or [u/v]' = (u'v-v'u)/v^2

OpenStudy (anonymous):

(fog)'=g'(x)f'(g(x))

OpenStudy (anonymous):

how would u differentiate (1+2x)^2

OpenStudy (anonymous):

wait, i think i got it

OpenStudy (anonymous):

by the way we are done with the solution, if you simplify the fraction you would get the required answer

OpenStudy (anonymous):

so, 1/2 ln(3u+6) => g(x)=3u+6 f(x)=1/2ln(g(x))

OpenStudy (anonymous):

omg, yes. thank you so much!

OpenStudy (anonymous):

you're more than welcome, this was good practicing .... all the best